Differential Equations, please find the general solution

2 5. (t) = - ( 11 1 21 1| c(t) 2 11/
Answer #1

x() = 2. 27 x (치 2 let n () = 씨 호 3) then - () (*) 72. 2. (t) 버 0 2.. 키지 (서게 (치 +의치 4275 (서 키 (D- 12 -2 22} = 0 (1) 지42길 473& ns =20,tų + x3 7 (1) +12) (0-1) = 22 n 2 3 10²-30 +1) x₂ - (20-3) 2=0 7 (oriaz = 2c₂ett- + (0-1) 33 = 0 :: –(3) -M, -23 + ((1) is, The ausiliary earn is, (m + 1) = 0 complementary solu Partictan solum **- tong +36 Cett (2 - 1 1 5 ) - Cze- teet wher